OpenStudy (anonymous):

Express the answers to the following calculations in scientific notation (a) 0.0095 + (8.5 x 10^-3) (b) 653 + (5.75 x 10^-8) (c) 850,000 – (9.0 x 10^5) (d) (3.6 x 10^-4) x (3.6 x 10^6) Please show work, thank you :D

OpenStudy (anonymous):

first you got to get common exponents so because these are fairly small numbers i would just convert your 8.5x10^-3 into reg notation and do the problem then change it back

OpenStudy (anonymous):

the first one can be done directly as they are both 10^-3 9.5 x 10^-3 + 8.5 x 10^-3 = 18 x 10^-3

OpenStudy (anonymous):

the others are best done as zbay suggests

OpenStudy (anonymous):

(b) convert (5.75 x 10^-8) to reg: = 0.0000000575 + 635 =635.0000000575 = 6.350000000575 x 10^2

OpenStudy (anonymous):

(c) = 8.5 x 10^5 - 9.0 x 10^5 = -0.5 x 10^5 = -5 x 10^4 (d) 3.6^2 x 10^2 = 12.96 x 10^2 = 1.296 x 10^3